OpenStudy (anonymous):

What Are the solutions of 3x^2-11x-4=0 ?

OpenStudy (anonymous):

use the quadratic formula: \(\large x=\frac{-b\pm \sqrt{b^2-4ac}}{2a} \) where a=3, b=-11, c=-4.

OpenStudy (anonymous):

check that... factoring is easier... can you factor that expression?

OpenStudy (anonymous):

(3x+1)(x-4)=0 x=4 or -1/3

OpenStudy (anonymous):

If someone can factor it I can do it

OpenStudy (anonymous):

Thanks @shelbyconig

OpenStudy (anonymous):

x = (-(-11) +/- sqrt((-11)^2 - 4(3)(-4))) / (2*3) x = (11 +/- sqrt(121 + 48)) / 6 x = (11 +/- sqrt(169)) / 6 x = (11 +/- 13) / 6 x = (24 or -2) / 6 x = 4 or -1/3 if you need more here you go,, and your welcome!